Silven Industries, which manufactures and sells a highly successful line of summer lotions and insect repellents, has decided to diversify in order to stabilize sales throughout the year. A natural area for the company to consider is the production of winter lotions and creams to prevent dry and chapped skin.

         After considerable research, a winter products line has been developed. However, Silven�s president has decided to introduce only one of the new products for this coming winter. If the product is a success, further expansion in future years will be initiated.

         The product selected (called Chap-Off) is a lip balm that will be sold in a lipstick-type tube. The product will be sold to wholesalers in boxes of 24 tubes for $7 per box. Because of excess capacity, no additional fixed manufacturing overhead costs will be incurred to produce the product. However, a $50,000 charge for fixed manufacturing overhead will be absorbed by the product under the company�s absorption costing system.

         Using the estimated sales and production of 100,000 boxes of Chap-Off, the Accounting Department has developed the following cost per box:


  Direct materials $ 3.30
  Direct labor 1.80
  Manufacturing overhead 1.20
  Total cost $ 6.30

The costs above include costs for producing both the lip balm and the tube that contains it. As an alternative to making the tubes, Silven has approached a supplier to discuss the possibility of purchasing the tubes for Chap-Off. The purchase price of the empty tubes from the supplier would be $1.30 per box of 24 tubes. If Silven Industries accepts the purchase proposal, direct labor and variable manufacturing overhead costs per box of Chap-Off would be reduced by 10% and direct materials costs would be reduced by 25%.

  

Required:
1a.

Calculate the total variable cost of producing one box of Chap-Off? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  

1b.

Assume that the tubes for the Chap-Off are purchased from the outside supplier, calculate the total variable cost of producing one box of Chap-Off? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  

1c. Should Silven Industries make or buy the tubes?
Make
Buy

   

2.

What would be the maximum purchase price acceptable to Silven Industries? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

     

3.

Instead of sales of 100,000 boxes, revised estimates show a sales volume of 130,000 boxes. At this new volume, additional equipment must be acquired to manufacture the tubes at an annual rental of $42,000. Assume that the outside supplier will not accept an order for less than 130,000 boxes.


a.

Calculate the total relevant cost of making 130,000 boxes and total relevant cost of buying 130,000 boxes. (Do not round intermediate calculations.)

I NEED AN ANSWER TO THE LAST QUESTION UNDERLINED ABOVE. I TRIED ALOT BUT CAN'T FIND IT/...PLZ HELP!!
Answer #1
Concepts and reason

Variable Cost: A variable cost refers to the cost which is directly related to per unit of goods manufactured. The variable rises as the volume of production increases.

Sales: Sale of any goods or services can be made on cash or credit basis. The amount receivable on sale can either be received immediately in cash or such a payment can be received at some future date. In case of sale is being made on credit basis the company maintains an account of such customer in its books as Debtor or Accounts Receivable.

Cost: Cost is the amount expended to any particular product or amount related to any particular product.

Fundamentals

Direct materials: These are the materials which are used in the Production process which constitutes an integral portion of the final product, the cost of which is identifiable and traceable directly to it. The cost of direct material can be directly chargeable to the final product as compared to the indirect material.

Direct labor: Workers or employees directly involved in the manufacture of goods and services are called the direct labor. Direct labor can be directly traceable to a specific product or cost center. In other words, the cost of the workers which is directly involved in the transformation of the raw material into the finished product.

Overheads: It is the indirect expenses of the business. Overheads are of two types (1) Fixed overheads and (2) Variable overheads.

Fixed overheads: these are the expenses which do not change with the change of production. These expenses are fixed up to a certain level of production after that level the fixed expenses are increased.

1a.

Calculate the variable manufacturing overheads per box using the equation as shown below:

Variablemanufacturingoverheads=((Totalproduction×Manufacturingcost)Fixedmanufacturingoverheads)Totalproduction=(100,000×$1.20)$50,000100,000=$120,000$50,000100,000=$70,000100,000\begin{array}{c}\\{\rm{Variable manufacturing overheads}} = \frac{{\left( \begin{array}{l}\\\left( {{\rm{Total production}} \times {\rm{Manufacturing cost}}} \right) - \\\\{\rm{Fixed manufacturing overheads}}\\\end{array} \right)}}{{{\rm{Total production}}}}\\\\ = \frac{{\left( {{\rm{100,000}} \times {\rm{\$ 1}}{\rm{.20}}} \right) - {\rm{\$ 50,000}}}}{{{\rm{100,000}}}}\\\\ = \frac{{{\rm{\$ 120,000}} - {\rm{\$ 50,000}}}}{{{\rm{100,000}}}}\\\\ = \frac{{{\rm{\$ 70,000}}}}{{{\rm{100,000}}}}\\\end{array}

=$0.70 = {\rm{\$ 0}}{\rm{.70}}

Hence, the variable manufacturing overheads are $.070.

Calculate the total variable cost per box using the equation as follows:

Totalvariablecost=Directmaterial+Directlabor+Variablemanufacturingoverheads=$3.30+$1.80+$0.70=$5.80\begin{array}{c}\\{\rm{Total variable cost}} = {\rm{Direct material}} + {\rm{Direct labor}} + {\rm{Variable manufacturing overheads}}\\\\ = {\rm{\$ 3}}{\rm{.30}} + {\rm{\$ 1}}{\rm{.80}} + {\rm{\$ 0}}{\rm{.70}}\\\\ = {\rm{\$ 5}}{\rm{.80}}\\\end{array}

1b.

Calculate the total variable cost per box using the equation as shown below:

Totalvariablecost=(Directmaterial+Purchasecostoftube+Directlaborcost+Variablemanufacturingoverhead)=$2.48+$1.30+$1.62+$.63=$6.03\begin{array}{c}\\{\rm{Total variable cost}} = \left( \begin{array}{l}\\{\rm{Direct material}} + {\rm{Purchase cost of tube}} + {\rm{Direct labor cost}}\\\\ + {\rm{Variable manufacturing overhead}}\\\end{array} \right)\\\\ = {\rm{\$ 2}}{\rm{.48}} + {\rm{\$ 1}}{\rm{.30}} + {\rm{\$ 1}}{\rm{.62}} + {\rm{\$ }}{\rm{.63}}\\\\ = {\rm{\$ 6}}{\rm{.03}}\\\end{array}

Working notes:

(i)

Calculate the direct material cost using the equation as follows:

Directmaterial=Directmaterialcost×(1Reductioninmaterialcost)=$3.30×(10.25)=$3.30×0.75=$2.48\begin{array}{c}\\{\rm{Direct material}} = {\rm{Direct material cost}} \times \left( {{\rm{1}} - {\rm{Reduction in material cost}}} \right)\\\\ = {\rm{\$ 3}}{\rm{.30}} \times \left( {{\rm{1}} - {\rm{0}}{\rm{.25}}} \right)\\\\ = {\rm{\$ 3}}{\rm{.30}} \times {\rm{0}}{\rm{.75}}\\\\ = {\rm{\$ 2}}{\rm{.48}}\\\end{array}

Hence, the direct material cost is $2.48.

(ii)

Calculate the direct labor cost using the equation as follows:

Directlaborcost=Laborcostperunit×(1Reductioninlaborcost)=$1.80×(1.10)=$1.80×0.90=$1.62\begin{array}{c}\\{\rm{Direct labor cost}} = {\rm{Labor cost per unit}} \times \left( {{\rm{1}} - {\rm{Reduction in labor cost}}} \right)\\\\ = {\rm{\$ 1}}{\rm{.80}} \times \left( {{\rm{1}} - {\rm{.10}}} \right)\\\\ = {\rm{\$ 1}}{\rm{.80}} \times {\rm{0}}{\rm{.90}}\\\\ = {\rm{\$ 1}}{\rm{.62}}\\\end{array}

Hence, the direct labor cost is $1.62.

(iii)

Calculate the variable manufacturing overhead using the equation as follows:

Variablemanufacturingoverhead=(Variablemanufacturingoverhead×(1Reductioninmanufacturingcost))=($.7×(10.10))=$.70×.90=$.63\begin{array}{c}\\{\rm{Variable manufacturing overhead}} = \left( \begin{array}{l}\\{\rm{Variable manufacturing overhead}} \times \\\\\left( {{\rm{1}} - {\rm{Reduction in manufacturing cost}}} \right)\\\end{array} \right)\\\\ = \left( {{\rm{\$ }}{\rm{.7}} \times \left( {{\rm{1}} - {\rm{0}}{\rm{.10}}} \right)} \right)\\\\ = {\rm{\$ }}{\rm{.70}} \times {\rm{.90}}\\\\ = {\rm{\$ }}{\rm{.63}}\\\end{array}

Hence, the variable manufacturing overhead is $.63.

1c.

The total variable cost, if the tubes are manufactured is $5.80 and the total variable cost, if the tubes are purchased, is $6.03. The variable cost of manufacturing the tubes is less costly in comparison to the variable cost of purchasing the tubes.

2.

Calculate the maximum purchase price using the equation as shown below:

Maximumpurchaseprice=(Purchasecostoftube+Variablecost(iftubesismanufactured)Variablecost(iftubesispurchased))=$1.30+$5.80$6.03=$1.07\begin{array}{c}\\{\rm{Maximum purchase price}} = \left( \begin{array}{l}\\{\rm{Purchase cost of tube}} + {\rm{Variable cost}}\\\\\left( {{\rm{if tubes is manufactured}}} \right)\\\\ - {\rm{Variable cost }}\left( {{\rm{if tubes is purchased}}} \right)\\\end{array} \right)\\\\ = {\rm{\$ 1}}{\rm{.30}} + {\rm{\$ 5}}{\rm{.80}} - {\rm{\$ 6}}{\rm{.03}}\\\\ = {\rm{\$ 1}}{\rm{.07}}\\\end{array}

3a.

Calculate the total relevant cost of making 130,000 boxes using the equation as shown below:

Totalrelevantcost=(Salesvolume×Variablemanufacturingcost)+Rentalofmachine=(130,000×$5.80)+$42,000=$754,000+$42,000=$796,000\begin{array}{c}\\{\rm{Total relevant cost}} = \left( {{\rm{Sales volume}} \times {\rm{Variable manufacturing cost}}} \right)\\\\ + {\rm{Rental of machine}}\\\\ = \left( {{\rm{130,000}} \times {\rm{\$ 5}}{\rm{.80}}} \right) + {\rm{\$ 42,000}}\\\\ = {\rm{\$ 754,000}} + {\rm{\$ 42,000}}\\\\ = {\rm{\$ 796,000}}\\\end{array}

Calculate the total relevant cost of buying 130,000 boxes using the equation as shown below:

Totalrelevantcost=Salesvolume×Variablebuyingcost=130,000×$6.03=$783,900\begin{array}{c}\\{\rm{Total relevant cost}} = {\rm{Sales volume}} \times {\rm{Variable buying cost}}\\\\ = {\rm{130,000}} \times {\rm{\$ 6}}{\rm{.03}}\\\\ = {\rm{\$ 783,900}}\\\end{array}

Ans:

The total variable cost per box is $5.80.

The total variable cost per box is $6.03.

S industries should make the tubes.

The maximum purchase price is $1.07.

The total relevant cost of making 130,000 boxes is $796,000.

The total relevant cost of buying 130,000 boxes is $783,900.